FTIR OBSERVATION OF C_{n} AND Si_{m}C_{n} CLUSTERS FORMED IN Ar MATRICES BY LASER ABLATION

\underline{SHUNLIN WANG}, XIANDE DING, and W.R.M. GRAHAM, Department of Physics, Texas Christian University, Fort Worth, TX 76129.

Well resolved vibrational spectra of C_{n} and Si_{m}C_{n} clusters have been produced by trapping products from the pulsed laser evaporation of pure carbon and silicon/carbon rods in Ar at 10 K. Controlled thermally induced diffusion in the matrix has enabled the growth of large clusters. To assist in the analysis of vibrational spectra extensive carbon-13 isotopic shift measurements have been made. Objectives in this work are the identification of linear C_{n} clusters with n ¿ 9 or cyclic C_{n} c lusters, and Si_{m}C_{n} clusters consisting of long carbon chains with terminal Si atoms.